# Towards Understanding the Benefits of Neural Network Parameterizations in Geophysical Inversions: A Study With Neural Fields

### v1

@@ Abstract
Recent research in test-time machine learning methods has shown that some machine learning models, without any prior learning, can improve the results of geophysical inversions. Some examples include the Deep Image Prior Inversions (DIP-Inv) and the Neural Fields Inversions (NFs-Inv), where the inverse problems are reparametrized by the weights of the machine learning models. In this work, we employ neural fields, which use neural networks to map a coordinate to the corresponding physical property value at that coordinate, in a test-time learning manner. For a test-time learning method, the weights are learned during the inversion, as compared to traditional approaches which require a network to be trained using a training data set. Results for synthetic examples in seismic tomography and direct current resistivity inversions are shown first. We then perform a singular value decomposition analysis on the Jacobian of the weights of the neural network (SVD analysis) for both cases to explore the effects of neural networks on the recovered model. The results show that the test-time learning approach can eliminate unwanted artifacts in the recovered subsurface physical property model caused by the sensitivity of the survey and physics. Therefore, NFs-Inv improves the inversion results compared to the conventional inversion in some cases such as the recovery of the dip angle or the prediction of the boundaries of the main target. In the SVD analysis, we observe similar patterns in the left-singular vectors as were observed in some diffusion models, trained in a supervised manner, for generative tasks in computer vision. This observation provides evidence that there is an implicit bias, which is inherent in neural network structures, that is useful in supervised learning and test-time learning models. This implicit bias has the potential to be useful for recovering models in geophysical inversions.
@@ I Introduction
Geophysical inversion is an important step for interpreting observed geophysical measurements [1], [2]. In an inversion, a physical property model of the subsurface is estimated from measurements obtained on the surface, in boreholes or from airborne systems. The quality of the inversion results depends on robust forward modelling, reliable prior information, and well-designed regularization techniques. Due to the inherent ill-posed nature of geophysical inverse problems, developing a suitable regularization is challenging. Conventional regularization typically penalizes spatial variations and/or explicitly promotes proximity to a reference model [3], but with advances in machine learning, there are opportunities to explore more advanced regularization.
@@
Recently, machine learning methods have been demonstrated to be successful in many geophysical applications. Alyousuf et al. [4] trained ML models to identify unexpected ${\rm CO_{2}}$ migration. Zhang et al. [5] improved the reconstructions of subsurface seismic velocities by mapping velocity models to raw seismic waveforms using trained conditional adversarial networks. Liu et al. [6] predicted anomalous bodies in the subsurface by matching apparent resistivity data to the resistivity model using trained conventional neural networks.
@@
Neural fields are advanced ML methods that can solve problems in 3D rendering, point cloud reconstruction, biomedical image recovery, etc [7], [8], [9], [10]. Neural fields use coordinate-based neural networks to map spatial, temporal, or other coordinates to the corresponding quantity at those coordinates. Neural fields are also called coordinate-based signal representations, or neural implicit representations in certain contexts. Unlike the previously described data-driven methods, which map input-output pairs, neural fields use, and view, neural networks as a continuous representation of the target’s properties in the coordinate space. Many applications of neural fields are not data-driven, instead, they are physics-informed or test-time learning.
@@
A significant and successful application of neural fields is in 3D rendering, which aims to generate new views of complex scenes using a set of input views [8]. The output values (colors and opacity) of the neural fields are fed into a forward map, that is based on a volume rendering algorithm, to generate the predicted observations (pixel values). The objective function, which measures the difference between the predicted measurements and real measurements, is used to update the weights of the neural network so that the loss value is expected to decrease iteratively. This training process is very similar to the process of solving inverse problems, so neural fields have been adapted to many inverse problems since then. The benefits of utilizing neural fields come from the following factors: (1) There is no need for pre-training, namely, the performance of these methods is not limited by the quality and quantity of the training dataset. (2) In principle, neural fields are mesh-free methods, namely, we can employ an adaptive grid scheme during inversion, which may bring benefits in terms of the accuracy of the reconstruction and the computational costs. (3) Parameterizing the inverse problems in a continuous setting naturally introduces smoothing regularization effects. (4) We can leverage the implicit regularization/bias from the neural network’s structure and optimization method.
@@
The performance of neural fields methods in 3D rendering has been shown to be much better than that of conventional methods in terms of distinguishing the boundaries between objects, predicting the pixel values of the main objects, etc [8]. The sources of the benefits of utilizing the neural fields are still controversial. For example, Yu et al. [11] argue that the neural network is not the source of these benefits by showing that when the neural network in the NeRF, a typical NFs method in 3D rendering, is replaced with spherical harmonics, the performance in the 3D rendering tasks is not affected. Many other works do not explicitly discuss the sources of benefits by employing the neural fields methods in 3D rendering [8], [12]. A key consideration when assessing and designing a model is its performance in representation tasks. For example, in the context of 3D rendering, researchers want to find a model that can reconstruct the 2D image better by learning self-similarity. The main strategies for improving model performance include: selecting a positional encoding function tailored to the problem [13], [14], improving the activation function [15], and employing different bases [16], [11]. Notably, the majority of the works still utilize neural networks as a component in the neural field methods partly thanks to the well-established programming frameworks, such as Tensorflow [17] and PyTorch [18], that simplify the implementation of the neural networks.
@@
Neural fields methods have also garnered significant attention from researchers outside of the domain of computer vision and computer graphics. A main domain for testing the usage of neural field methods is biomedical imaging [19], [20], [10], [21], [22]. In CT and MRI imaging reconstruction tasks, Shen et al. [21] showed that learning the implicit neural representation of the reconstructed images by adapting a prior embedding can significantly improve the resolution of the reconstruction results. Liu et al. [22] used neural fields with a pre-trained denoising regularizer and showed that this approach could improve the recovery of 3D refractive index images from discrete measurements. Sun et al. [10] improve reconstruction performance in CT tomographic imaging with a coordinate-based neural network by generating new measurements from existing data and incorporating these new measurements during training.
@@
In the geosciences, Rasht-Behesht et al. [23] proposed physics-informed neural networks for seismic inversions by parameterizing two physical quantities in the wave equation with two coordinate-based neural networks. The results show that this method can provide a good recovery of the subsurface model without the need for an explicit regularization term. Sun et al. [24] tested the performance of parameterizing the seismic velocity directly with a coordinate-based neural network in the full-waveform inversion. Both works argue that neural fields methods can improve the reconstruction performance in scenarios lacking accurate starting models. Neural fields have also been applied to the joint inversion of seismic and electrical data. Liu et al. [25] used a Bayesian neural network to parameterize the global model properties and a convolutional neural network to parameterize the reservoir properties. The numerical results show that the proposed pipeline provides a more accurate recovery of the reservoir model. Notably, some advanced techniques in computer vision and computer graphics were adapted to some of the scientific inverse problems mentioned above and proved to be useful. For example, both Rasht-Behesht et al. [23] and Sun et al. [24] mentioned the importance of using an adaptive sinusoidal activation function. Liu et al. [22] showed the benefits of adopting a task-specific positional encoding function. There are only a limited number of works in geophysical inverse problems that use neural fields methods, and a majority of them focus on seismic. Thus, there is substantial potential to further explore this new area.
@@
It is well-known that learning alone is insufficient to explain the good performance of deep networks in terms of generalizability. The fact that deep neural networks can describe the underlying distribution of a dataset, even in high dimensional spaces that suffer from the curse of dimensionality, suggests that there is a prior, or implicit bias, that has been imposed in the training process [26]. However, the underlying mechanism of this implicit bias has not been fully explained. For example, some researchers have found that certain modern ML models’ structures (e.g. Transformers, CNNs, DNNs) have an implicit regularization or bias that is beneficial for some inverse problems [27], [28], [29], [30]. Some of them suggested that the benefits come from the high dimensionality of the neural network [31], [32], [33], [30]. Others analyze this generalizability using probabilistic theory [34], [35]. Although some of these works have taken a mathematical point of view and have aimed to prove rigorous theorems, they have been limited to overly simplified settings. Other works used the empirical-based method (or the so-called “scientific method”), which forms hypotheses and designs controlled experiments to test them [36]. Specifically, it highlights the importance of empirical analyses of deep neural networks to validate or refute existing theories and assumptions, as well as to address fundamental questions regarding the mechanisms behind their successes. Although relatively under-explored, these empirical approaches offer significant potential to strengthen our understanding of deep learning and to drive substantial advancements in both theoretical and practical domains. Some well-known works that emphasize empirical-based methods include: Physics of Large Language Models [36], where Zhu et al. used synthetic data and built an idealized environment for the training of LLMs to find the “laws of all LLMs”. A scaling law for LLMs states that the performance of LLMs scales with model size, dataset size, and the amount of compute power [37]. Kadkhodaie et al. [26] find that in the training of diffusion models (a class of generative models), the model tends to learn a basis that is adapted to geometric features in the underlying images, and they name these bases as “geometry-adaptive harmonic bases”.
@@
Kadkhodaie et al. [26] attribute the implicit bias to the model structure and optimization method, not the training data. Therefore, we can utilize the analysis of that work to take one step further to examine the performance of the implicit bias and test-time learning methods in the geophysical inverse problems. Previous works have demonstrated the benefits of test-time learning methods to geophysical inversions and have attributed these benefits to the implicit bias provided by the machine learning model structures [38], [39]. However, these works do not provide an analysis of what prior(s) this implicit bias imposes during the inversion. In this work, we will show that test-time machine learning methods can improve the geophysical inversion result by finding weights that can capture geometric structures in the physical property model.
@@
In this work, we will first introduce our NFs-Inv method and its applications in seismic tomography and direct current resistivity (DCR) inversions. To avoid ambiguity, we use the same structure of the Neural Network (NN) for both cases. However, we use different positional encoding functions in different cases since some positional encoding functions are helpful in terms of learning high-frequency variations of the subsurface model. The comparisons between the conventional Tikhonov-style inversion results and the NFs-Inv inversion results clearly show that our NFs-Inv method can improve the recovery of the subsurface model in some cases. We will then present our analysis of these inversion results. Our analysis, which uses a singular value decomposition (SVD) analysis on the Jacobian of the weights of the neural network, will partly explain why searching over the NN-weights space will improve the inversion results. To the best of our knowledge, this is the first study that applies neural field methods to seismic tomography inversions and direct current resistivity inversions and gives evidence to show that the benefits of the TTL method are from the implicit bias inherited from the NN’s structure.
@@
    Presenting its application on linear/non-linear geophysical inversions with homogeneous/inhomogeneous background. We compare the NFs-Inv and conventional inversion results when no explicit regularization is used. These results demonstrate that the implicit bias from the NFs-Inv is appropriate for some geophysical inversion scenarios.
@@
    Providing an empirically-based explanation of the benefits of employing NFs-Inv. Our analysis of the Jacobian shows similar structures to the “geometry-adaptive harmonic representations” described by Kadkhodaie et al. [26]. These structures enable the inversion to overcome the problems of predicting unwanted artifacts due to sensitivity, such as near-electrode artifacts or the concentration along the ray path. This result suggests that the implicit bias in the NFs-Inv could be similar to the implicit bias found in certain dataset-trained machine learning models, such as diffusion models.
@@ II Methodology
Many geophysical inversion methods can be categorized into a physics-driven approach or a data-driven approach. The physics-driven approach imposes explicit regularization to the objective function or uses a problem-specific constrained algorithm during the inversion process. The advantage is that it is a white-boxed approach–namely, the regularization and constraints have well-established geological, mathematical and/or statistical theories as a foundation. The physics-driven approach can be utilized when there is only limited knowledge of the survey domain, but it sometimes can be very time-consuming since it needs to run a forward simulation in each epoch during the test time. The data-driven approach typically requires more prior knowledge of the geological structures in the subsurface within the survey domain. This prior knowledge can be used to build a dataset of plausible subsurface geophysical models and then can be learned by an ML model trained using this dataset. The trained ML models can give fast predictions for new observations. This approach is black-boxed—namely, it is not guaranteed that the trained ML model can return a plausible subsurface model that fits the observations for any test data. Our work will combine elements of the two approaches: The proposed method is physics-informed since it includes a forward numerical simulation, which will make the method more robust and guarantee that the inversion results can fit the observations to a desired level. The proposed method also directly utilizes the implicit regularization from the coordinate-based representations without the need for a training dataset.
@@ II-A Background and conventional inversion approach
For cross-hole seismic tomography, the goal of the inversion is to recover the subsurface velocity model by analyzing the travel times of seismic waves between boreholes. In this survey, a seismic source is placed in one borehole, and seismic receivers (geophones) are positioned in another borehole. The seismic source generates waves that travel through the subsurface and are detected by the receivers in the neighbouring borehole. By moving the sources and receivers to various depths within their respective boreholes, multiple ray paths are recorded. The travel times of these waves are then used to create a seismic velocity model of the subsurface, which is related to the lithology and elastic properties of the subsurface. Using the straight-ray assumption, we can calculate the arrival time at each receiver for a given source using equation 1.
@@
where $\sigma$ is the electrical conductivity (which is the reciprocal of electrical resistivity $\rho=1/\sigma$), $\phi$ is the electric potential, and $\mathbf{j}_{source}$ is the source current density. We solve this equation using a finite volume method where we discretize the subsurface. For the examples we will show, we use a 2D tensor mesh. The mesh consists of a core region, with uniform cells, and padding regions on the lateral boundaries and the bottom of the mesh with expanding cells to ensure boundary conditions are satisfied. The forward modelling is implemented using SimPEG [43].
@@
The objective function is a summation of the data misfit term $\phi_{d}(m,d^{obs})$ and the regularization term $\phi_{m}(m)$ with a trade-off parameter that weights the relative combination of each term. We assume that the noise is Gaussian and depends on the magnitude of the voltage measurements in the DCR survey, so an uncertainties matrix $W_{d}$ has been used to quantify the noise [2]. For the cross-hole seismic tomography survey, we assume that the noise is Gaussian but is independent of the magnitude of the arrival time, so we adapt an identity matrix for $W_{d}$.
@@
Often the norms applied are $p_{s,x,z}=2$, however, norms of less than 2 can also be employed to promote sparse or compact structures [3]. Regularization terms are commonly employed for the cases where we have no or limited geological prior knowledge. More advanced regularization can also be employed. For example, Astic and Oldenburg [44] proposed an inversion framework for integrating the petrophysical and geological prior into the direct current resistivity inversion. Laloy et al. [45] and Lopez-Alvis et al. [46] integrated the prior of distribution of the spatial patterns of the subsurface materials into the inverse process using a trained Variational Autoencoders. In this work, we will perform conventional inversions using the regularization in equation 4 as our “benchmark” scenarios.

### v2

@@ Abstract
Recent research in test-time machine learning methods has shown that some machine learning models, without any prior learning, can improve the results of geophysical inversions. Some examples include the Deep Image Prior Inversions (DIP-Inv) and the Neural Fields Inversions (NFs-Inv), where the inverse problems are reparametrized by the weights of the machine learning models. In this work, we employ neural fields, which use neural networks to map a coordinate to the corresponding physical property value at that coordinate, in a test-time learning manner. For a test-time learning method, the weights are learned during the inversion, as compared to traditional approaches, which require a network to be trained using a training dataset. Results for synthetic examples in seismic tomography and direct current resistivity inversions are shown first. We then perform a singular value decomposition analysis on the Jacobian of the weights of the neural network (SVD analysis) for both cases to explore the effects of neural networks on the recovered model. The results show that the test-time learning approach can eliminate unwanted artifacts in the recovered subsurface physical property model caused by the sensitivity of the survey and physics. Therefore, NFs-Inv improves the inversion results compared to the conventional inversion in some cases, such as the recovery of the dip angle or the prediction of the boundaries of the main target. In the SVD analysis, we observe similar patterns in the left-singular vectors as were observed in some diffusion models, trained in a supervised manner, for generative tasks in computer vision. This observation provides evidence that there is an implicit bias, which is inherent in the neural network structures, that is useful in supervised learning and test-time learning models. This implicit bias has the potential to be useful for recovering models in geophysical inversions.
@@ I Introduction
Geophysical inversion is an important step for interpreting observed geophysical measurements [1], [2], [3]. In an inversion, a physical property model of the subsurface is estimated from measurements obtained on the surface, in boreholes or from airborne systems. The quality of the inversion results depends on robust forward modelling, reliable prior information, and well-designed regularization techniques. Due to the inherently ill-posed nature of geophysical inverse problems, developing a suitable regularization is challenging. Conventional regularization typically penalizes spatial variations and/or explicitly promotes proximity to a reference model [4], [5], but with advances in machine learning, there are opportunities to explore more advanced regularization.
@@
Recently, machine learning methods have been demonstrated to be successful in many geophysical applications [6], [7], [8], [9]. Alyousuf et al. [7] trained ML models to identify unexpected ${\rm CO_{2}}$ migration. Zhang et al. [8] improved the reconstructions of subsurface seismic velocities by mapping velocity models to raw seismic waveforms using trained conditional adversarial networks. Liu et al. [9] predicted anomalous bodies in the subsurface by matching apparent resistivity data to the resistivity model using trained conventional neural networks.
@@
Neural fields are advanced ML methods that can solve problems in 3D rendering, point cloud reconstruction, biomedical image recovery, etc [10], [11], [12], [13], [14]. Neural fields use coordinate-based neural networks to map spatial, temporal, or other coordinates to the corresponding quantity at those coordinates. Neural fields are also called coordinate-based signal representations, or neural implicit representations, in certain contexts. Unlike the previously described data-driven methods, which map input-output pairs, neural fields use and view neural networks as a continuous representation of the target’s properties in the coordinate space. Many applications of neural fields are not data-driven, instead, they are physics-informed or test-time learning.
@@
A significant and successful application of neural fields is in 3D rendering, which aims to generate new views of complex scenes using a set of input views [11]. The output values (colors and opacity) of the neural fields are fed into a forward map, which is based on a volume rendering algorithm, to generate the predicted observations (pixel values). The objective function, which measures the difference between the predicted measurements and real measurements, is used to update the weights of the neural network so that the loss value is expected to decrease iteratively. This training process is very similar to the process of solving inverse problems, so neural fields have been adapted to many inverse problems since then. The benefits of utilizing neural fields come from the following factors: (1) There is no need for pre-training, namely, the performance of these methods is not limited by the quality and quantity of the training dataset. (2) In principle, neural fields are mesh-free methods, namely, we can employ an adaptive grid scheme during inversion, which may bring benefits in terms of the accuracy of the reconstruction and the computational costs. (3) Parameterizing the inverse problems in a continuous setting naturally introduces smoothing regularization effects. (4) We can leverage the implicit/inductive regularization/bias from the neural network’s structure and optimization method.
@@
The performance of neural fields methods in 3D rendering has been shown to be much better than that of conventional methods in terms of distinguishing the boundaries between objects, predicting the pixel values of the main objects, etc [11]. The sources of the benefits of utilizing the neural fields are still controversial. For example, Yu et al. [15] argue that the neural network is not the source of these benefits by showing that when the neural network in the NeRF, a typical NFs method in 3D rendering, is replaced with spherical harmonics, the performance in the 3D rendering tasks is not affected. Many other works do not explicitly discuss the sources of benefits by employing the neural fields methods in 3D rendering [11], [16]. A key consideration when assessing and designing a model is its performance in representation tasks. For example, in the context of 3D rendering, researchers want to find a model that can reconstruct the 2D image better by learning self-similarity. The main strategies for improving model performance include: selecting a positional encoding function tailored to the problem [17], [18], improving the activation function [19], and employing different bases [20], [15]. Notably, the majority of the works still utilize neural networks as a component in the neural field methods.
@@
Neural fields methods have also garnered significant attention from researchers outside of the domain of computer vision and computer graphics. A main domain for testing the usage of neural fields methods is biomedical imaging [21], [22], [13], [23], [24]. In CT and MRI imaging reconstruction tasks, Shen et al. [23] showed that learning the implicit neural representation of the reconstructed images by adapting a prior embedding can significantly improve the resolution of the reconstruction results. Liu et al. [24] used neural fields with a pre-trained denoising regularizer and showed that this approach could improve the recovery of 3D refractive index images from discrete measurements. Sun et al. [13] improve reconstruction performance in CT tomographic imaging with a coordinate-based neural network by generating new measurements from existing data and incorporating these new measurements during training.
@@
In the geosciences, Rasht-Behesht et al. [25] proposed physics-informed neural networks for seismic inversions by parameterizing two physical quantities in the wave equation with two coordinate-based neural networks. The results show that this method can provide a good recovery of the subsurface model without the need for an explicit regularization term. Sun et al. [26] tested the performance of parameterizing the seismic velocity directly with a coordinate-based neural network in the full-waveform inversion. Both works argue that neural fields methods can improve the reconstruction performance in scenarios lacking accurate starting models. Neural fields have also been applied to the joint inversion of seismic and electrical data. Liu et al. [27] used a Bayesian neural network to parameterize the global model properties and a convolutional neural network to parameterize the reservoir properties. The numerical results show that the proposed pipeline provides a more accurate recovery of the reservoir model. Notably, some advanced techniques in computer vision and computer graphics were adapted to some of the scientific inverse problems mentioned above and proved to be useful. For example, both Rasht-Behesht et al. [25] and Sun et al. [26] mentioned the importance of using an adaptive sinusoidal activation function. Liu et al. [24] showed the benefits of adopting a task-specific positional encoding function. There are some works in geophysical inverse problems on parameterizing the subsurface models using neural networks [28], [29], but there are only a limited number of works that use neural fields methods, and a majority of them focus on seismic inversions. Thus, there is substantial potential to further explore this new area.
There are many previous works on Physics-Informed Neural Networks (PINNs), which also use coordinate-based neural networks in solving scientific inverse problems. In general, PINNs incorporate governing physical laws, such as partial differential equations, directly into deep neural networks (DNNs) as prior knowledge. This approach addresses the challenge of data scarcity in training ML models to solve scientific inverse problems [30], [31], [32]. Compared to the direct-feedforward networks, PINNs can achieve better accuracy and shorter training time by using a smaller training dataset [33], [34]. Our work does not fall into the category of PINN mainly because we do not let the neural network learn the physics of the problem. Namely, we enforce the PDE constraints by using a PDE solver to calculate the residual instead of using a PDE residual where the PDE is directly embedded in the objective function, so there is no need for any training dataset.
@@
It is well-known that learning alone is insufficient to explain the good performance of deep networks in terms of generalizability. The fact that deep neural networks can describe the underlying distribution of a dataset, even in high-dimensional spaces that suffer from the curse of dimensionality, suggests that there is a prior (implicit/inductive bias) that has been imposed in the training process [35]. However, the underlying mechanism of this implicit bias has not been fully explained. For example, some researchers have found that certain modern ML models’ structures (e.g. Transformers, CNNs, DNNs) have an implicit regularization or bias that is beneficial for some inverse problems [36], [37], [38], [39]. Some of them suggested that the benefits come from the high dimensionality of the neural network [39], [40], [41]. Others analyze this generalizability using probabilistic theory [42], [43]. Although some of these works have taken a mathematical point of view and have aimed to prove rigorous theorems, they have been limited to overly simplified settings. Other works used the empirical-based method (or the so-called “scientific method”), which forms hypotheses and designs controlled experiments to test them [44]. Specifically, it highlights the importance of empirical analyses of deep neural networks to validate or refute existing theories and assumptions, as well as to address fundamental questions regarding the mechanisms behind their successes. Although relatively under-explored, these empirical approaches offer significant potential to strengthen our understanding of deep learning and to drive substantial advancements in both theoretical and practical domains. Some well-known works that emphasize empirical-based methods include: Physics of Large Language Models [44], where Zhu et al. used synthetic data and built an idealized environment for the training of LLMs to find the “laws of all LLMs”. A scaling law for LLMs states that the performance of LLMs scales with model size, dataset size, and the amount of compute power [45]. Kadkhodaie et al. [35] found that the diffusion models tend to learn a basis that is adapted to geometric features in the underlying images, and they name these bases as “geometry-adaptive harmonic bases”.
@@
Kadkhodaie et al. [35] attributed the implicit bias to the model structure and optimization method, not the training data alone. Therefore, we can utilize the analysis of that work to take one step further to examine the performance of the implicit bias and test-time learning methods in the geophysical inverse problems. Previous works have demonstrated the benefits of test-time learning methods to geophysical inversions and have attributed these benefits to the implicit bias provided by the machine learning model structures [46], [47]. However, these works do not provide an analysis of what priors this implicit bias impose during the inversion. In this work, we will show that test-time machine learning methods can improve the geophysical inversion result by finding weights that can capture geometric structures in the physical property model.
@@
In this work, we will first introduce our NFs-Inv method and its applications in seismic tomography and direct current resistivity (DCR) inversions. We use the same structure of the Neural Network (NN) for both cases. However, we use different positional encoding functions in different cases since some positional encoding functions are helpful in terms of learning high-frequency variations of the subsurface model. The comparisons between the conventional Tikhonov-style inversion results and the NFs-Inv inversion results clearly show that our NFs-Inv method can improve the recovery of the subsurface model in some cases. We will then present our analysis of these inversion results. Our analysis, which uses a singular value decomposition (SVD) analysis on the Jacobian of the weights of the neural network, will partly explain why searching over the NN-weights space will improve the inversion results. To the best of our knowledge, this is the first study that applies neural fields methods to seismic tomography inversions and direct current resistivity inversions and gives evidence to show that the benefits of the test-time learning (TTL) method are from the implicit bias inherited from the NN’s structure.
@@
    Presenting its application on linear/non-linear geophysical inversions with homogeneous/inhomogeneous background. We compare the NFs-Inv and conventional inversion results. These results demonstrate that the implicit bias from the NFs-Inv is appropriate for some geophysical inversion scenarios.
@@
    Providing an empirically-based explanation of the benefits of employing NFs-Inv. The left-singular vectors of the Jacobian have a similar pattern to the “geometry-adaptive harmonic bases” described by Kadkhodaie et al. [35]. The basis of the Jacobian reveals what has been imposed by the implicit/inductive bias. The bias enables the inversion to overcome the problems of predicting unwanted artifacts due to sensitivity, such as near-electrode artifacts or the concentration along the ray path. This result suggests that the implicit bias in the NFs-Inv could be similar to the implicit bias found in certain dataset-trained machine learning models, such as diffusion models.
@@ II Methodology
Many geophysical inversion methods can be categorized into a physics-driven approach or a data-driven approach. The physics-driven approach imposes explicit regularization to the objective function or uses a problem-specific constrained algorithm during the inversion process. The advantage is that it is a white-boxed approach–namely, the regularization and constraints have well-established geological, mathematical and/or statistical theories as a foundation. The physics-driven approach can be utilized when there is only limited knowledge of the survey domain, but it can sometimes be very time-consuming since it needs to run a forward simulation in each epoch during the test time. The data-driven approach typically requires more prior knowledge of the geological structures in the subsurface within the survey domain. This prior knowledge can be used to build a dataset of plausible subsurface geophysical models and then can be learned by the ML model trained using this dataset. The trained ML models can give fast predictions for new observations. This approach is black-boxed—namely, it is not guaranteed that the trained ML model can return a plausible subsurface model that fits the observations for any test data. Our work will combine elements of the two approaches: The proposed method is physics-informed since it includes a forward numerical simulation, which will make the method more robust and guarantee that the inversion results can fit the observations to a desired level. The proposed method also directly utilizes the implicit regularization from the coordinate-based representations without the need for a training dataset.
@@ II-A Background and conventional inversion approach
For cross-hole seismic tomography, the goal of the inversion is to recover the subsurface velocity model by analyzing the travel times of seismic waves between boreholes. In this survey, a seismic source is placed in one borehole, and seismic receivers (geophones) are placed in another borehole. The seismic source generates waves that travel through the subsurface and are detected by the receivers in the neighboring borehole. By moving the sources and receivers to various depths within their respective boreholes, multiple arrival times are recorded. The travel times of these waves are then used to create a seismic velocity model of the subsurface, which is related to the lithology and elastic properties of the subsurface. Using the straight-ray assumption, we can calculate the arrival time at each receiver for a given source using equation 1.
@@
where $\sigma$ is the electrical conductivity (which is the reciprocal of electrical resistivity $\rho=1/\sigma$), $\phi$ is the electric potential, and $\mathbf{j}_{source}$ is the source current density. We solve this equation using a finite volume method. For the examples we will show, we use a 2D tensor mesh. The mesh consists of a core region, with uniform cells, and padding regions on the lateral boundaries and the bottom of the mesh with expanding cells to ensure boundary conditions are satisfied. The forward modelling is implemented using SimPEG [51].
@@
The objective function is a summation of the data misfit term $\phi_{d}(m,d^{obs})$ and the regularization term $\phi_{m}(m)$ with a trade-off parameter that weights the relative combination of each term. We assume that the noise is Gaussian and depends on the magnitude of the voltage measurements in the DCR survey. A matrix $W_{d}$, which includes data uncertainties, has been used to quantify the noise [2]. For the cross-hole seismic tomography survey, we assume that the noise is Gaussian but is independent of the magnitude of the arrival time, so we adapt an identity matrix for $W_{d}$.
@@
Often, the norms applied are $p_{s,x,z}=2$, however, norms of less than 2 can also be employed to promote sparse or compact structures [4]. Regularization terms are commonly employed for the cases where we have no or limited geological prior knowledge. More advanced regularization can also be employed. For example, Astic and Oldenburg [52] proposed an inversion framework for integrating the petrophysical and geological prior into the direct current resistivity inversion. Laloy et al. [53] and Lopez-Alvis et al. [54] integrated the prior of distribution of the spatial patterns of the subsurface materials into the inverse process using a trained Variational Autoencoder. In this work, we will perform conventional inversions using the regularization in equation 4 as the benchmark. The hyperparameter values in equation 4 can be found in Section III.